OWG Meeting 25-May-2011

From GlueXWiki
Jump to: navigation, search


  • Announcements
  • Review of minutes from 27-Apr-2011 meeting
  • Coil test update - Elliott, Yi, Beni, Ivan
    • short to ground, ground fault detector
    • strain gauges
    • VME fast DAQ system
    • SiPM tests
    • VME/VXS and HV/LV crate tests
  • IT review - Mark, Dave
  • Remote Access review - Graham
  • IRMIS - Elliott
  • Event size and raw event format - Dave, Carl
  • Event display - Elliott, Andrew
  • Java coda object framework - Elliott, Andrew
  • Other hardware tests - Chris, Beni
    • Multi-crate trigger test
    • FDC cosmic ray test
  • DAQ Group code distribution, build system, include dirs - Carl
  • Controls test stand - Hovanes, Elliott
  • Update on conditions database - Mark
  • Electronics, trigger and DAQ status reports - Chris, Alex, Dave, Fernando, Ben, Ed, William, Bryan
  • Mantis Task Tracker


1:30 PM Wed 25-May-2011 CC F326


Next Meeting

1:30 PM Wed 8-Jun-2011 CC F326


Present: Elliott W, Simon T, Bryan M, Graham H, Dave L, Hovanes E, Beni Z, Dave A.

Coil Test

Elliott reported on the status of the recent coil test.

  • Ramped to 1500 A the previous Mon.
  • No problems with cryo, magnet, power supply, control system.
  • Short to ground requried addendum to OSP, caused a few days delay. Can run with max one short to ground.
  • A few problems with safety system.
    • some strain gauges broken or improperly wired.
    • needed another OSP revision to allow operation with no strain gauges on radial Heim columns.
    • used field asymmetry measurments to indirectly determine stresses on radial Heim columns.
  • VME fast DAQ worked fine.
  • SiPM tests went well.
  • VME/VXS crate tests went well.
    • some fans locked up at around 180 G, expect max 70 G in Hall D.
  • Coil 2 put back together, found one leak that will be repaired.

IT Review

Dave reported on a few recommendations from the review concerning offsite requirements, authority within the Offline group, need more JLab employees if we are to perform 40% of the coding effort, need for better justification of numbers presented in the review. There will eventually be an external review later this year.

Remote Access review

Graham reported on the ongoing review and planning for remote access to controls in the hall.

  • Read-only access is no problem. EPICS gateways will allow anyone to view EPICS variables.
  • Two-factor authentication will be required to control devices in the hall, about $100 per authentication device.
  • Questions about remote shift taking, running DAQ vs controlling hardware.


Elliott reported that IRMIS running under Glassfish is working. Scot, Mark and Elliott will test it with information from the coil test controls system.

Event Size

Dave reported that recent simulations show the BCAL data may be 1 kB larger than expected due to noise and changes in readout segmentation.

Event Display and Coda Objects

Elliott reported that he will meet with the CNU students this week.

Trigger Test

Bryan reported that modules are still coming in. They have tested two FADC250's in a single crate and all tests went well.

Controls Test Stand

Hovanes reported that the racks are in but there is still some facilities work to be done in EEL 126.

Mantis Task Tracker

A few new tasks have been entered, more to go.